I need help on solving this equation?
(2x)/(x+3)+5=(3)/(x+3)



Answer by Mathtut(3670)

2x+5(x+3)=3.........multiplied all terms by x+3
:
2x+5x+15=3
:
7x=-12
:
x=-12/7